OpenStudy (anonymous):

-5 and 1/3 x 1 and 2/5 Answer Choices 1)-7 and 7/15 2) 6 and 4/5 3) -7 and 2/15 4) 6 and 1/3

OpenStudy (jdoe0001):

\(\bf -5\cfrac{1}{3}\times 1\cfrac{2}{5}\\ \quad \\ \quad \\ \textit{recall that}\quad a\cfrac{b}{c}\implies \cfrac{a\times c+b}{c}\quad \textit{and that}\quad \cfrac{a}{b}\times \cfrac{c}{d}\implies \cfrac{a\times b}{c\times d}\)